# A Bayesian Approach to Segmentation with Noisy Labels via Spatially Correlated Distributions

### v1

@@ 3.1 Overview
$=\sum_{i}\mathbb{E}_{q(y_{i}^{*}|y_{i},\eta_{i})q(\eta_{i})}\left[\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]$
$-\sum_{i}\mathbb{E}_{q(y_{i}^{*}|y_{i},\eta_{i})q(\eta_{i})}\left[\log q(y_{i}^{*}|y_{i},\eta_{i})\right]-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right]$
$+\sum_{i}\mathbb{E}_{q(y^{*}_{i}|y_{i},\eta_{i})q(\eta_{i})}\left[\log p(y_{i}|y^{*}_{i},\eta_{i})\right],$ (3)
where $q(\eta_{i})$ is the marginal distribution of $q(\bm{\eta})$, which can be easily obtained when $q(\bm{\eta})$ follows a Gaussian distribution $q(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{m},\bm{\Gamma})$. $q(y_{ik}^{*}|y_{ic},\eta_{i})$ and $p(y_{ik}|y_{ic}^{*},\eta_{i})$ are defined as $q(y_{ik}^{*}|y_{ic},\eta_{i})=\delta_{kc}(1-r(\eta_{i}))+(1-\delta_{kc})r(\eta_{i})V_{kc}$, $p(y_{ik}|y_{ic}^{*},\eta_{i})=\delta_{kc}(1-r(\eta_{i}))+(1-\delta_{kc})r(\eta_{i})W_{kc}$, respectively. Here, $\delta_{kc}$ is the Kronecker delta, $r(\cdot)$ is the sigmoid function, and $V\in[0,1]^{{|\mathcal{Y}|}\times{|\mathcal{Y}|}},\ \sum_{c=1}^{|\mathcal{Y}|}V_{kc}=1,\ V_{kk}=0\ (k,c\in\mathcal{Y})$ represents the probability of the clean label when the given noisy label $c$ is inconsistent with the clean label $k$, and $W\in[0,1]^{{|\mathcal{Y}|}\times{|\mathcal{Y}|}},\ \sum_{c=1}^{|\mathcal{Y}|}W_{kc}=1,\ W_{kk}=0\ (k,c\in\mathcal{Y})$ represents the probability of the noisy label when the given clean label $c$ is inconsistent with the noisy label $k$.
Owing to the assumption of conditional independence, the expectations in the first, third, and fifth terms in the ELBO (3.1) are tractable, as we can sum over the possible labels for each pixel independently. The spatial correlation of label errors is incorporated through the second and fourth terms in the ELBO (3.1). These terms involve the trace and inverse of large matrices $\bm{\Sigma}$ and $\bm{\Gamma}$, making direct computation infeasible. However, these computational challenges are efficiently addressed by leveraging the properties of KMS-structured covariance.
@@ 3.2 KMS-structured covariance
and the determinant is computed as $|R_{\rho}|=(1-\rho^{2})^{n-1}.$ See Appendix A.1 for mathematical derivations.
@@
This matrix is useful for representing correlations among one-dimensional variables, where the correlation decays exponentially as the index distance between two variables increases regardless of their absolute positions. Suppose a one-dimensional random variable $\mathbf{x}$ has a covariance structure given by: $\mathbb{E}[(x_{i}-\mu_{i})(x_{j}-\mu_{j})]=\sigma_{i}\sigma_{j}\rho^{|i-j|}$, where $\mu_{i}$ is the mean of $x_{i}$. By utilizing the KMS matrix $R_{\rho}$, the covariance matrix $\bm{\Sigma}$ can be expressed as $\bm{\Sigma}=VR_{\rho}V$ where $V$ is a diagonal matrix whose $(i,i)$-th element is $\sigma_{i}$. Consequently, the inverse of the covariance matrix is given by $\bm{\Sigma}^{-1}=V^{-1}R_{\rho}^{-1}V^{-1}$.
In the case of two-dimensional spatial correlations, such as those in images, we utilize the Kronecker product. Suppose a two-dimensional random variable $\mathbf{x}$ has a covariance structure given by $\mathbb{E}[(x_{ij}-\mu_{ij})(x_{uv}-\mu_{uv})]=\sigma_{ij}\sigma_{uv}\rho^{|i-u|+|j-v|}$ where $\mu_{ij}$ is the mean of $x_{ij}$. Its matrix form $\bm{\Sigma}$ can be expressed as $\bm{\Sigma}=V(R_{\rho}\otimes R^{\prime}_{\rho})V$. Here, we assume that the two-dimensional pixel index $(i,j)$ can be appropriately mapped to a one-dimensional index $n$. $V$ is a diagonal matrix whose $(n,n)$-th element corresponds $\sigma_{ij}$, and $\otimes$ denotes the Kronecker product. Due to the properties of KMS matrices and the Kronecker product, the inverse covariance matrix is given by $\bm{\Sigma}^{-1}=V^{-1}(R_{\rho}^{-1}\otimes R^{\prime-1}_{\rho})V^{-1}$. Since $V$ is a diagonal matrix and the nonzero off-diagonal elements in the inverses of the KMS matrices $R_{\rho}$ and $R^{\prime}_{\rho}$ exist only in adjacent rows and columns, the quadratic form in the exponent of the Gaussian distribution $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})\propto\exp(-\frac{1}{2}(\bm{\eta}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{\eta}-\bm{\mu}))$ involves only pairs of adjacent pixel variables. This structure makes the conditional distribution Markovian, meaning that, $p(\eta_{i}|\bm{\eta}_{\backslash i})=p(\eta_{i},\bm{\eta}_{\backslash i})/p(\bm{\eta}_{\backslash i})\propto\exp(-\sum_{j\in{\mathcal{N}}_{i}}\Sigma^{-1}_{ij}(\eta_{i}-\mu_{i})(\eta_{j}-\mu_{j}))$, which leads to $p(\eta_{i}|\bm{\eta}_{\backslash i})=p(\eta_{i}|\{\eta_{j}|j\in{\mathcal{N}}_{i}\})$ where $\bm{\eta}_{\backslash i}$ denotes all variables in $\bm{\eta}$ except $\eta_{i}$, and ${\mathcal{N}}_{i}$ represents the set of adjacent pixels of $i$. In other words, the KMS-structured covariance matrix ensures that the Gaussian distribution retains a Markov Random Field-like property while maintaining the tractability of both the inverse and the determinant of the covariance matrix. The dependence of this conditional distribution on $\rho$ is illustrated in Figure 1(b). As shown in Figure 1(b), the model can confidently infer label errors when adjacent pixels are also likely to contain label errors and the spatial correlation is strong ($\rho$ is high), even if such errors are rare.
@@ 3.3 Optimization
For simplicity, we define $N$ as the number of training samples, $\bm{\omega}$, as the prior distribution parameters, $\{\bm{\nu}^{(n)}\}_{n=1}^{N}$, as the image-specific posterior parameters, and $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}^{(n)})$ as the negative ELBO (3.1) computed using these parameters. Our training algorithm is outlined in Algorithm 1, where we illustrate the case with a single sample per step, though this can be generalized to multiple samples. In each step, the model and prior parameters are updated $K$ times, while the posterior parameters are updated $M$ times. After training, the optimized model parameters are used to predict clean labels for new input images. Note that in Algorithm 1, the prior parameters $\bm{\omega}$ remain fixed, but with a sufficiently large training dataset, they can also be optimized.
Algorithm 1 Training loop of the model and posterior parameters optimization.
1:$\bm{\theta},\bm{\omega},\{\bm{\nu}_{0}^{(\ell)}\}_{\ell=1}^{N}\leftarrow$ Initialize parameters
@@
5:         Compute $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}_{0}^{(n)})$ and its gradient with respect to $\bm{\theta}$
@@
9:         Compute $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}_{m}^{(n)})$ and its gradient with respect to $\bm{\nu}_{m}^{(n)}$
@@ 4.1 Dataset
It is common practice in noisy-label segmentation research to artificially introduce label noise for evaluation because most publicly available datasets come with clean annotations [20, 31, 39]. Following this convention, we also employ synthetic noise in our experiments, with details of our noise generation process described in Section 4.2.1.
JSRT Dataset [30] is a publicly available dataset provided by the Japanese Society of Radiological Technology (JSRT), which comprises chest radiographs annotated with segmentation labels for the lung fields, heart, and clavicles. All images are of size $256\times 256$ pixels, with 199 images designated for training and 50 for testing. Following the assumption common in noisy-label segmentation research—that no clean validation data is available—we evaluated our approach using the published split of 199 training images and 50 test images. Following a previous study [20], the Clavicle region was cropped to a fixed $96\times 224$ area for training and inference.
WHU Building Dataset [13] is a publicly available dataset for building segmentation containing satellite images of Christchurch with a resolution of 0.075 m covering an area of $450\,\text{km}^{2}$. Manual annotations were provided for $22{,}000$ buildings. The satellite images were downsampled to a resolution of 0.3 m and then divided into $512\times 512$ tiles. Among the non-overlapping tiles, $4{,}736$ were used for training, 1,036 for validation, and 2,416 for testing. Similar to JSRT, we adopted the published split for the training and testing data.
@@ 4.2 Experimental Setup
Figure 2: Examples of Noisy Labels. In medical imaging, label noise is often simulated using morphological operations such as erosion, dilation, and affine transformations [20]. Conversely, building segmentation commonly involves omission noise, commission errors, and boundary noise [31]. The pink region indicates labels present in the noisy but absent in the clean label, while the blue region indicates the opposite.
In the JSRT dataset, we evaluated the impact of noise caused by boundary error, which is a critical issue in the medical imaging domain [39, 43]. In accordance with previous works [37, 44, 37, 20], we employed morphological transformations as synthetic noise. As illustrated in Figure 2, the noise settings consider not only boundary ambiguities (i.e., dilation or erosion) but also affine transformations. We examined the effect of noise using $\alpha$ o represent the proportion of affected data and $\beta$ to determine its strength, setting $\{(\alpha,\beta)\}=\{(0.3,0.5),(0.5,0.7),(0.7,0.7)\}$.
In the WHU Building dataset, we modeled three types of label noise: omission noise, where labels are missing in regions that should be labeled [31, 21]; commission error, where labels are present in regions that should not be due to temporal changes [31]; and boundary noise, which commonly occurs in segmentation tasks [31], as illustrated in Figure 2. We define the omission ratio ($\phi$), the commission ratio ($\zeta$), and the boundary noise probability ($\lambda$) as the proportions of the total instances in the image that are perturbed by noise. Specifically, the noise parameters are set as $\{(\phi,\zeta,\lambda)\}=\{(0.1,0.0,0.3),\ (0.2,0.05,0.5),\ (0.3,0.1,0.7)\}$.
@@
We compare our method against several baselines, including cross-entropy loss (CE), noisy-label learning approaches for classification tasks (CoT [10], JoCoR [34], EM [1]), noise-robust losses for segmentation T-Loss [9], as well as techniques that account for partial spatial correlations, namely Superpixel (SP) [20] and SpatialCorrect (SC) [39]. For each method, we adopted the hyperparameters reported in the publicly available implementations or original papers, assuming them to be the best practices for their respective approaches. Although proposed method supports multi-label segmentation, many of the baseline methods are designed for binary segmentation. Therefore, in JSRT dataset, we conducted experiments by treating each class (Lung, Heart, Clavicle) as a separate binary segmentation task.
@@
We use a U-Net [28] with an EfficientNet-B0 [32] encoder, pre-trained on ImageNet [4]. The model is optimized using Adam [18] with a learning rate of 0.001 and batch sizes of 16 (JSRT) and 32 (WHU Building). The parameters in the prior $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})$ and the initial values of the parameters in the variational posterior $q(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{m},\bm{\Gamma})$ are set as follows. The parameters $\bm{\mu}$, the diagonal elements of $\bm{\Sigma}$, initial values of the elements of $\bm{m}$ and the diagonal elements of $\bm{\Gamma}$ are all same and denoted as $\mu$, $\sigma^{2}$, $m$, and $\gamma^{2}$, respectively. These parameters were tuned using $15\%$ of the WHU Building training set and validation sets, then fixed for all experiments, including JSRT, as follows: $\rho=0.75$, $m=-5$, $\gamma=1$, $\mu=-2$, and $\sigma=1$.

### v3

@@ 4.1 A Tractable Variational Objective via the ECCD
$=\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|y_{i},\eta_{i})}\left[\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})\right]-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]$
$-\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|y_{i},\eta_{i})}\left[\log q(y_{i}^{*}|y_{i},\eta_{i})\right]+\sum_{i}\mathbb{E}_{q(\eta_{i})q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log p(y_{i}|y^{*}_{i},\eta_{i})\right].$ (7)
A key advantage of this formulation is that the expectations over the discrete latent labels $\mathbf{y}^{*}$ in Eq.(4.1) are now decomposed into a sum over individual pixels. This avoids the intractable summation over all possible label map configurations that hindered the direct optimization of the marginal likelihood in Eq.(3), rendering the objective tractable. The overall goal is to optimize the model parameters $\bm{\theta}$ and the variational parameters of $q(\bm{\eta})$ and $q(y_{i}^{*}|y_{i},\eta_{i})$ to maximize this lower bound. The ELBO in Eq.(4.1) has an important interpretation as a principled generalization of the standard cross-entropy loss. To see this, we can analyze its components.
The first term in this ELBO (4.1) can be interpreted as an expected reconstruction term. By expanding the expectation over the discrete clean label $y_{i}^{*}$, it becomes a cross-entropy loss with soft labels:
$\sum_{i}\mathbb{E}_{q(y_{i}^{*}|y_{i},\eta_{i})q(\eta_{i})}\left[\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})\right]=\sum_{i=1}^{HW}\sum_{k=1}^{|\mathcal{Y}|}q(y_{ik}^{*}=1|y_{i})\log p_{\theta}(y_{ik}^{*}=1|\mathbf{x}),$ (8)
where $q(y_{ik}^{*}=1|y_{i})=\mathbb{E}_{q(\eta_{i})}[q(y_{ik}^{*}=1|y_{i},\eta_{i})]$ is the marginal posterior probability that the clean label for pixel $i$ is class $k$. This replaces the hard, fixed one-hot observed labels with soft, inferred posterior probabilities. Building on this interpretation of the first term, we can now demonstrate more formally that our entire objective is a principled generalization of the standard cross-entropy loss. This is shown by examining the conditions under which the remaining terms in the ELBO vanish.
The second and third terms together constitute the negative Kullback-Leibler (KL) divergence, $\text{KL}[q(\bm{\eta})||p(\bm{\eta})]$. This term vanishes when the variational posterior is identical to the prior, i.e., $q(\bm{\eta})=p(\bm{\eta})$. Similarly, the fourth and fifth terms represent the expected log-likelihood of the observed labels under the noise model, regularized by the entropy of the label posterior. These terms also cancel each other out under the assumption of a symmetric noise process, where $V=W$ in Eqs. (6) and (5). This condition is met in non-trivial noisy scenarios, for instance, with a uniform noise model where an error from any class is equally likely to transition to any other incorrect class (i.e., $W_{kc}=V_{kc}=1/(|\mathcal{Y}|-1)$ for $k\neq c$).
In such a setting, maximizing the ELBO is equivalent to maximizing only the first term. As we saw, if we assume a noise-free setting, the posterior distribution $q(y_{i}^{*}|y_{i},\eta_{i})$ collapses to a point mass (i.e., a Kronecker delta function centered at the observed label $y_{i}$), making the soft labels $q(y_{ik}^{*}|y_{i})$ identical to the observed one-hot labels. Consequently, our objective function reduces precisely to the standard cross-entropy loss. Therefore, our method provides a principled generalization of the standard supervised learning framework, where the additional terms in the ELBO serve to regularize the posterior and enforce spatial correlation in the inferred label errors.
While the expectations over the discrete labels $y_{i}^{*}$ are tractable, the terms involving the continuous variable $\bm{\eta}$ pose a significant computational challenge. The number of elements in $\bm{\eta}$ equals the number of image pixels ($HW$), which can be in the hundreds of thousands, and these elements are spatially correlated. For instance, the negative KL divergence term, $\text{KL}[q(\bm{\eta})||p(\bm{\eta})]$, is analytically computable since both $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})$ and $q(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{m},\bm{\Gamma})$ are Gaussians. Its value is given by:
$\text{KL}[q(\bm{\eta})||p(\bm{\eta})]=\frac{1}{2}\left(\log\frac{|\bm{\Gamma}|}{|\bm{\Sigma}|}+HW-{\rm{Tr}}\left(\bm{\Sigma}^{-1}\bm{\Gamma}\right)-(\bm{m}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{m}-\bm{\mu})\right).$ (9)
However, evaluating this term is computationally prohibitive for large images, as it requires computing the determinants and inverses of the $HW\times HW$ covariance matrices, in addition to the trace and quadratic form terms. These challenges are efficiently addressed by leveraging the properties of the KMS-structured covariance, as we describe next.
@@ 4.2 KMS-structured covariance
and the determinant is computed as $|R_{\rho}|=(1-\rho^{2})^{n-1}.$ See Appendix A.1 for mathematical derivations. While the KMS matrix is well-established in fields like signal processing for modeling AR(1) processes Grenander & Szegö (1958); Brockwell & Davis (1991), its unique properties — a sparse inverse and an analytic determinant — have been largely untapped as a computational tool in modern machine learning. Our work bridges this gap by demonstrating its utility as the key component for making the ECCD framework computationally tractable.
@@
This matrix is useful for representing correlations among one-dimensional variables, where the correlation decays exponentially as the index distance between two variables increases regardless of their absolute positions.
Suppose a one-dimensional random variable $\mathbf{x}$ has a covariance structure given by
$\mathbb{E}[(x_{i}-\mu_{i})(x_{j}-\mu_{j})]=\sigma_{i}\sigma_{j}\rho^{|i-j|},$ (10)
where $\mu_{i}$ is the mean of $x_{i}$. By utilizing the KMS matrix $R_{\rho}$, the covariance matrix can be expressed as
$\bm{\Sigma}=VR_{\rho}V,\quad\bm{\Sigma}^{-1}=V^{-1}R_{\rho}^{-1}V^{-1},$ (11)
where $V$ is a diagonal matrix whose $(i,i)$-th element is $\sigma_{i}$.
In the case of two-dimensional spatial correlations, such as those in images, we utilize the Kronecker product. Suppose a two-dimensional random variable $\mathbf{x}$ has a covariance structure given by
$\mathbb{E}[(x_{ij}-\mu_{ij})(x_{uv}-\mu_{uv})]=\sigma_{ij}\sigma_{uv}\rho^{|i-u|+|j-v|},$ (12)
where $\mu_{ij}$ is the mean of $x_{ij}$. Its matrix form can be expressed as
$\bm{\Sigma}=V(R_{\rho}\otimes R^{\prime}_{\rho})V,\quad\bm{\Sigma}^{-1}=V^{-1}(R_{\rho}^{-1}\otimes R_{\rho}^{\prime-1})V^{-1},$ (13)
where we assume that the two-dimensional pixel index $(i,j)$ is mapped to a one-dimensional index $n$. $V$ is a diagonal matrix whose $(n,n)$-th element corresponds to $\sigma_{ij}$, and $\otimes$ denotes the Kronecker product.
Since $V$ is diagonal and $R_{\rho}^{-1}$ is tridiagonal, the inverse covariance $\bm{\Sigma}^{-1}$ is a sparse matrix. The properties of the KMS matrix provide an elegant solution to the computational challenges outlined in Section 4.1. First, the determinant of the covariance matrix can be computed analytically, bypassing the need for expensive numerical methods. Second, the sparse structure of the inverse matrix is crucial for the remaining terms. It reduces the complexity of computing the quadratic form and the trace term in the KL divergence (9) from $O((HW)^{2})$ to $O(HW)$, which is linear in the number of pixels.
Beyond this computational speed-up, the sparse structure of the inverse also ensures that the quadratic form
$(\bm{\eta}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\bm{\eta}-\bm{\mu})$ (14)
[…]
In other words, the KMS-structured covariance ensures that the Gaussian distribution retains a Markov Random Field property while maintaining the tractability of both the inverse and the determinant of the covariance matrix. The effect of the correlation parameter $\rho$ is illustrated in Figure 1(b). As shown, the model can confidently infer label errors when adjacent pixels are also likely to contain errors and the spatial correlation is strong (i.e., $\rho$ is high).
@@ 4.3 Optimization
We formulate the optimization problem as follows. Let $N$ be the number of training samples. We denote the parameters of the prior distribution as $\bm{\omega}$, the image-specific parameters of the variational posterior as $\{\bm{\nu}^{(n)}\}_{n=1}^{N}$, and the negative ELBO for the $n$-th sample (from Eq.(4.1)) as $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}^{(n)})$.
Our training procedure, outlined in Algorithm 1, employs an alternating optimization scheme. For clarity, the algorithm is presented for a single sample per step, though this can be generalized to mini-batches. In each step for a given sample, the model parameters $\theta$ are updated $K$ times, while the corresponding variational posterior parameters $\bm{\nu}^{(n)}$ are updated $M$ times. After training, the optimized model parameters $\theta$ are used for inference on new images. Note that in Algorithm 1, the prior parameters $\bm{\omega}$ are treated as fixed,but they could also be optimized jointly with a sufficiently large training dataset.
Algorithm 1 Training loop for alternating optimization of model and posterior parameters.
[…]
@@
5:   Compute $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}_{0}^{(n)})$ and its gradient w.r.t. $\bm{\theta}$
@@
9:   Compute $\mathcal{L}(\bm{\theta},\bm{\omega},\bm{\nu}_{m}^{(n)})$ and its gradient w.r.t. $\bm{\nu}_{m}^{(n)}$
@@ 5.1 Dataset
It is common practice in noisy‐label segmentation research to artificially introduce label noise for evaluation because most publicly available datasets come with clean annotations Li et al. (2021); Song et al. (2022); Yao et al. (2023). Following this convention, we also employ synthetic noise in our experiments, with details of our noise generation process described in Section 5.2.1.
JSRT Dataset Shiraishi et al. (2000) is a publicly available dataset provided by the Japanese Society of Radiological Technology (JSRT), which comprises chest radiographs annotated with segmentation labels for the lung fields, heart, and clavicles. All images are of size $256\times 256$ pixels, with 199 images designated for training and 50 for testing. Following the assumption common in noisy-label segmentation research—that no clean validation data is available—we evaluated our approach using the published split of 199 training images and 50 test images. Following a previous study Li et al. (2021), the Clavicle region was cropped to a fixed $96\times 224$ area for training and inference.
WHU Building Dataset Ji et al. (2018) is a publicly available dataset for building segmentation containing satellite images of Christchurch with a resolution of 0.075 m covering an area of $450\,\text{km}^{2}$. Manual annotations were provided for 22,00022{,}000 buildings. The satellite images were downsampled to a resolution of 0.3 m and then divided into $512\times 512$ tiles. Among the non-overlapping tiles, 4,7364{,}736 were used for training, 1,036 for validation, and 2,416 for testing. Similar to JSRT, we adopted the published split for the training and testing data.
@@ 5.2 Experimental Setup
Figure 2: Examples of Noisy Labels. In medical imaging, label noise is often simulated using morphological operations such as erosion, dilation, and affine transformations Li et al. (2021). Conversely, building segmentation commonly involves omission noise, commission errors, and boundary noise Song et al. (2022). The pink region indicates labels present in the noisy but absent in the clean label, while the blue region indicates the opposite.
In the JSRT dataset, we evaluated the impact of noise caused by boundary error, which is a critical issue in the medical imaging domain Yao et al. (2023); Zhang et al. (2020b). In accordance with previous works Xue et al. (2020); Zhang et al. (2020c); Xue et al. (2020); Li et al. (2021), we employed morphological transformations as synthetic noise. As illustrated in Figure 2, the noise settings consider not only boundary ambiguities (i.e., dilation or erosion) but also affine transformations. We examined the effect of noise using $\alpha$ to represent the proportion of affected data and $\beta$ to determine its strength, setting $\{(\alpha,\beta)\}=\{(0.3,0.5),(0.5,0.7),(0.7,0.7)\}$.
In the WHU Building dataset, we modeled three types of label noise: omission noise, where labels are missing in regions that should be labeled Song et al. (2022); Liu et al. (2024); commission error, where labels are present in regions that should not be due to temporal changes Song et al. (2022); and boundary noise, which commonly occurs in segmentation tasks Song et al. (2022), as illustrated in Figure 2. We define the omission ratio ($\phi$), the commission ratio ($\zeta$), and the boundary noise probability ($\lambda$) as the proportions of the total instances in the image that are perturbed by noise. Specifically, the noise parameters are set as $\{(\phi,\zeta,\lambda)\}=\{(0.1,0.0,0.3),\ (0.2,0.05,0.5),\ (0.3,0.1,0.7)\}$.
@@
We compare our method against several baselines, including cross-entropy loss (CE), noisy-label learning approaches for classification tasks (CoT Han et al. (2018), JoCoR Wei et al. (2020), EM Bekker & Goldberger (2016)), noise-robust losses for segmentation T-Loss Gonzalez-Jimenez et al. (2025), as well as techniques that account for partial spatial correlations, namely Superpixel (SP) Li et al. (2021) and SpatialCorrect (SC) Yao et al. (2023). For each method, we adopted the hyperparameters reported in the publicly available implementations or original papers, assuming them to be the best practices for their respective approaches. Although proposed method supports multi-label segmentation, many of the baseline methods are designed for binary segmentation. Therefore, in JSRT dataset, we conducted experiments by treating each class (Lung, Heart, Clavicle) as a separate binary segmentation task.
@@
We use a U-Net Ronneberger et al. (2015) with an EfficientNet-B0 Tan & Le (2019) encoder, pre-trained on ImageNet Deng et al. (2009). The model is optimized using Adam Kingma & Ba (2014) with a learning rate of 0.001 and batch sizes of 16 (JSRT) and 32 (WHU Building). The parameters in the prior $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})$ and the initial values of the parameters in the variational posterior $q(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{m},\bm{\Gamma})$ are set as follows. The parameters $\bm{\mu}$, the diagonal elements of $\bm{\Sigma}$, initial values of the elements of $\bm{m}$ and the diagonal elements of $\bm{\Gamma}$ are all same and denoted as $\mu$, $\sigma^{2}$, $m$, and $\gamma^{2}$, respectively. These parameters were tuned using $15\%$ of the WHU Building training set and validation sets, then fixed for all experiments, including JSRT, as follows: $\rho=0.75$, m=5m=-5, $\gamma=1$, $\mu=-2$, and $\sigma=1$.